\begin{thebibliography}{9}
\bibitem{Af06} A. Aftalion, Vortices in Bose-Einstein condensates.  {\em Progress in Nonlinear Differential Equations and their Applications} {\bf 67}, Birkh\"auser,  2006.

\bibitem{AnMaSpar} 
P. Antonelli, D. Marahrens,  C. Sparber,
On the Cauchy problem for nonlinear Schr\"odinger equations with rotation.
\textit{Discrete Contin. Dyn. Syst.} \textbf{32} (2012), no. 3, 703--715.

%\bibitem{BaCD10} V. Banica, R. Carles, T. Duyckaerts, Minimal blow-up solutions to the mass-critical inhomogeneous NLS equation. Communications in Partial Differential Equations 36 (3), 2010, 487--531.


%\bibitem{BanVisc15} Valeria Banica, Nicola Visciglia, Scattering for NLS with a delta potential. Apr 13 2015. math.AP
%math.MP arXiv:1504.02640v1  abs. We prove $H^1$ scattering for defocusing NLS with a delta potential and mass-supercritical nonlinearity, hence extending in an inhomogeneous setting the classical 1-D scattering results first proved by Nakanishi in the translation invariant case.

%\bibitem{Bao} W. Bao, \href{http://www.math.nus.edu.sg/~bao/PS/mantova.pdf}{GPE article.}

%\bibitem{BaoCai13} W. Z. Bao and Y. Y. Cai, Mathematical theory and numerical methods for Bose-Einstein condensation, Kinetic and Related Models 6 (2013), 1--135.

%\bibitem{BaoBenCai12} Weizhu Bao, Naoufel Ben Abdallah, and Yongyong Cai, \href{https://doi.org/10.1137/110850451}{Gross-Pitaevskii-Poisson Equations for Dipolar Bose-Einstein Condensate with Anisotropic Confinement}.  SIAM Journal on Mathematical Analysis, 2012, Vol. 44, No. 3 : pp. 1713-1741. \crr

%\bibitem{BaoBCai12} W. Bao, N. Ben Abdallah and Y. Cai, Gross-Pitaevskii-Poisson equations for dipolar Bose-Einstein condensate with anisotropic confinement. SIAM J. Math. Anal. 44 (2012),  1713--1741.

%\bibitem{BenMeSW05} N. Ben Abdallah, F. M\'ehats, C. Schmeiser, and R. M. Weish\"aupl, The nonlinear Schr\"odinger  equation with a strongly anisotropic harmonic potential, SIAM J. Math. Anal. 37 (2005),  pp. 189--199.

\bibitem{BaoCai15} W. Bao, Y. Cai, %\href{https://doi.org/10.1137/140979241}
{Ground states and dynamics of spin-orbit-coupled Bose--Einstein condensates}. 
{\em SIAM J.  Appl. Math.} {\bf 75} (2015), no. 2, 492--517.

\bibitem{BaoWMar05} W. Bao, H. Wang,  P.  Markowich, Ground, symmetric and central vortex states in rotating 
Bose-Einstein condensates. {\em Comm. Math. Sci.} {\bf 3} (2005), 57--88. 


%\bibitem{Bao} T. Bartsch amd Z.-Q. Wang, Existence and multiplicity results for some superlinearelliptic problems on $\R^N$ , Comm. Partial Differential Equations 20 (1995), 1725--1741.

%\bibitem{BaoDuZh06} W. Bao, Q. Du,  Y. Zhang, Dynamics of rotating Bose-Einstein condensates and its efficient and accurate numerical computation. SIAM J. Appl. Math., 66(3):758--786, 2006. doi: 10.1137/050629392.

%\bibitem{Bao} W. Bao, \href{http://www.newton.ac.uk/programmes/HOP/Abstract5/bao.html}{Bao workshop}

%\bibitem{BaoJinM03} 5. W. Bao, S. Jin, and P. A. Markowich, Numerical study of time-splitting spectral discretizationsof nonlinear Schr\"odinger equations in the semiclassical regimes, SIAM J.  Sci. Comput. 25 (2003), no. 1, 27--64.

\bibitem{BW98}  J. Bourgain, W. Wang, Construction of blowup solutions for the nonlinear Schr\"odinger equation with critical nonlinearity. {\em Ann. Scer. Norm. Sup.  Pisa Cl. Sci}.  {\bf 25} (1997),  197--215.   %Issue: 1-2, page 197-215

\bibitem{BaHaHuZheng}
N. Basharat, H. Hajaiej, Y. Hu,  S. Zheng, Threshold for blowup and stability for nonlinear Schr\"odinger equation with rotation.  Preprint.

\bibitem{Car} R. Carles,  Critical nonlinear Schr\"odinger equations with and without harmonic potential,
\textit{Math. Models Methods Appl. Sci.} \textbf{12} (2002), 1513--1523.

%\bibitem{Car}  R. Carles, Geometric optics and instability for semi-classical Schr?odinger equations, preprint, arXiv:math.AP/0505468, 2005.

%\bibitem{Car11time} R. Carles, \href{https://hal.archives-ouvertes.fr/hal-00426530/en/}{Nonlinear Schr\"odinger equation with time dependent potential.}  Commun. Math. Sci. 9 (2011), no. 4, 937--964. 

%We prove a global well-posedness result for defocusing nonlinear Schrodinger equations with time dependent potential. We then focus on time dependent harmonic potentials. This aspect is motivated by Physics (Bose-Einstein condensation), and appears also as a preparation for the analysis of the propagation of wave packets in a nonlinear context. The main aspect considered here is the growth of high Sobolev norms of the solution.


%\bibitem{CarGa15} Remi Carles, Cl\'ement Gallo, \href{https://arxiv.org/abs/1502.00994}{Scattering for the nonlinear Schrodinger equation with a general one-dimensional confinement}. arXiv:1502.00994 [math.AP], 16pp.

%(3 Feb 2015)  We consider the defocusing nonlinear Schr{\"o}dinger equation in several space dimensions, in the presence of an external potential depending on only one space vari-able. This potential is bounded from below, and may grow arbitrarily fast at infinity. We prove existence and uniqueness in the associated Cauchy problem, in a suitable functional framework, as well as the existence of wave operators when the power of the nonlinearity is sufficiently large. Asymptotic completeness then stems from at least two approaches, which are briefly recalled.
%	16 pages   DOI:	10.1063/1.4932604

%\bibitem{CarNa04}  R. Carles and Y. Nakamura, Nonlinear Schr\"odinger equations with Stark potential,Hokkaido Math. J. 33 (2004), no. 3, 719--729.

%\bibitem{CarSi15} R. Carles, \href{https://hal.archives-ouvertes.fr/hal-00823573/en/}{Large time behavior in nonlinear Schr\"odinger equation with time dependent potential.}   With Jorge Drumond Silva. Commun. Math. Sci. 13 (2015), no. 2, 443--460.
	
	
	
%\bibitem{Cartan1899} \href{https://en.wikipedia.org/wiki/Differential_form}{\'Elie Cartan} (1899), Sur certaines expressions diff\'erentielles et le probl\`eme de Pfaff, Annales scientifiques de l'\'Ecole Normale Sup\'erieure: 239--332.

%\bibitem{Caz93} T. Cazenave, An introduction to nonlinear Schr\"odinger equations, volume 26 of Text. Met. Mat. Univ. Fed. Rio de Jan., 1993.
 %An Introduction to Nonlinear Schršdinger Equations, 
% (Textos de M\'etodos. Matem\'aticos, Universidade Federal do Rio de Janeiro, 26, third edition)

%characterization of Q via variational method using some functionals.\crr 

%\bibitem{Caz03} T. Cazenave, Semilinear Schr\"odinger equations. Courant Lecture Notes in Mathematics, vol. 10, New York University Courant Institute of Mathematical Sciences, New York, 2003.  

\bibitem{CazE88}  T. Cazenave, M. Esteban,
%\href{https://books.google.com/books?hl=en&lr=&id=vn8C0mwx_ZUC&oi=fnd&pg=PA155&dq=stability+magnetic+schrodinger+equation&ots=ByvdOJaGWz&sig=poiI103lTsWvsy9htIENepASo1k#v=onepage&q=stability%20magnetic%20schrodinger%20equation&f=false}
 {On the stability of stationary states for nonlinear Schr\"odinger equations with an external magnetic field}.  {\em Mat. Apl. Comput}. {\bf 7} (1988), 155--168.

%\bibitem{CazLion82} T. Cazenave, P.-L. Lions, Orbital stability of standing waves for some Schr\"odinger equations.  Comm. Math. Phys. {\bf 85} (1982), 549--561.

\bibitem{CazW90}  T. Cazenave, F. Weissler, The Cauchy problem for the critical nonlinear Schr\"odinger equation in $H^s$.
\textit{Nonlinear Anal.} \textbf{14} (1990), no. 10, 807--836.


\bibitem{ChMarS16} A.~Cheskidov, D.~Marahrens, C.~Sparber, %\href{https://arxiv.org/abs/1506.04706v3}
 {Global attractor for a Ginzburg-Landau type model of rotating Bose-Einstein condensates}.   %(Submitted on 15 Jun 2015 (v1), last revised 5 Aug 
{\em Dyn.  Partial Differ. Eqn}. {\bf 14} (2017),  5--32.
%arXiv. 2016. %(this version, v3))

%We study the long time behavior of solutions to a nonlinear partial differential equation arising in the description of trapped rotating Bose-Einstein  condensates. The equation can be seen as a hybrid between the well-known nonlinear Schr\"odinger/Gross-Pitaevskii equation and the Ginzburg-Landau equation. We prove existence and uniqueness of global in-time solutions in the physical energy space and establish the existence of a global attractor within the associated dynamics. We also obtain basic structural properties of the attractor and an estimate on its Hausdorff and fractal dimensions.
%Comments:	25 pages; some more typos fixed; additional references added
%Subjects:	Analysis of PDEs (math.AP); Quantum Gases (cond-mat.quant-gas); Dynamical Systems (math.DS)
%MSC classes:	35Q55, 35A01

\bibitem{De}
A. De Bouard,   Nonlinear Schr\"odinger equations with magnetic fields. 
\textit{Differential Integral Equations.} \textbf{4} (1991), no. 1, 73--88.

%\bibitem{Darw12a} Mohamad Darwich, Blowup for the Damped $L^2$-Critical Nonlinear Schr\"odinger Equation.\href{https://arxiv.org/abs/1101.2763}{Advances in Differential Equations}. Volume 17, Numbers 3-4 (2012), 337--367.

%(Submitted on 14 Jan 2011 (v1), last revised 3 Jul 2012 (this version, v3))

%We consider the Cauchy problem for the L2-critical damped nonlinear Schr\"odinger equation. We prove existence and stability of finite time blowup dynamics with the log-log blow-up speed for $\norm{\nabla u(t)}_2$.		arXiv:1101.2763 [math.AP]
% 	(or arXiv:1101.2763v3 [math.AP] for this version)

%\bibitem{Darw12b} Mohamad Darwich, \href{https://arxiv.org/abs/1206.6082v4}{On the $L^2$-critical nonlinear Schr\"odinger Equation with a nonlinear damping}    (Submitted on 26 Jun 2012 (v1), last revised 15 Jan 2013 

%We consider the Cauchy problem for the L2-critical nonlinear Schr\"{o}dinger equation with a nonlinear damping. According to the power of the damping term, we prove the global existence or the existence of finite time blowup dynamics with the log-log blow-up speed for $\norm{\nabla u(t)}_2$

%\bibitem{Fan15} C.~Fan, %\href{https://arxiv.org/abs/1510.00961}
%$\mathrm{Log}$-$\log$ blow up solutions blow up at exactly $m$ points, arXiv:1510.00961 [math.AP].

%\bibitem{Fra85} G. Fraiman, Asymptotic stability of manifold of self-similar solutions in self-focusing, Sov. Phys. JETP 61 (2) (1985) 228--233.

%\bibitem{FiMe01} G. Fibich and F. Merle. Self-focusing on bounded domains. Phys. D, 155(2001):132?58.

% T. Hmidi and S. Keraani. Blowup theory for the critical nonlinear Schr¬odinger  equations revisited. Int. Math. Res. Not., 46(2005):2815?828.

\bibitem{FiMerRa}  G. Fibich, F. Merle, P. Raphael,
Proof of a spectral property related to the singularity formation for the $L^2$ critical nonlinear Schr\"odinger equation. 
\textit{Phys. D} \textbf{220} (2006), 1--13.


\bibitem{Gross61} E.  Gross, Structure of a quantized vortex in boson systems. {\em Nuovo Cimento} {\bf 20} 
(1961), no.3, 454--477.


\bibitem{HaoHsiaoLi1}
C. Hao, L. Hsiao,  H. Li,
Global well posedness for the Gross-Pitaevskii equation with an angular momentum rotational term in three dimensions.
\textit{J. Math. Phys.} \textbf{48} (2007), no. 10, 102105.

\bibitem{HaoHsiaoLi2}
C. Hao, L. Hsiao,  H. Li,
Global well posedness for the Gross-Pitaevskii equation with an angular momentum rotational term.
\textit{Math. Methods Appl. Sci.} \textbf{31} (2008), 655--664.

%\bibitem{KeelTao}  M. Keel and T. Tao,  Endpoint Strichartz estimates, \textit{Amer. J. Math.} \textbf{120} (1998), 955--980.

\bibitem{Kita80} H. Kitada, %\href{https://www.researchgate.net/publication/29772185_On_a_construction_of_the_fundamental_solution_for_Schrodinger_equations}
 {On a construction of the fundamental solution for Schr\"odinger equations.} 
{\em J. Fac. Sci. Univ. Tokyo Sec. IA} {\bf 27} (1980), 193--226.

%In this paper, applying the skip method in Fujiwara (3) to the theory of multi-products of Fourier integral operators in Kitada and H.Kumano-go (6), we give a construction of the fundamental solution for the Cauchy problem of a pseudo-differential equation of Schrodinger's type. We regard this construction as a multi-product of Fourier integral operators, and investigate the pointwise convergence of the phase function and that of the symbol. Here we use neither the solution of the Hamilton-Jacobi equation in (6) nor the action of the classical orbit in (2),(4).

%\bibitem{KopLan95} Kopell, N.; Landman, M., Spatial structure of the focusing singularity of the nonlinear Schr\"odinger equation: a geometrical analysis. SIAM J. Appl. Math. {\bf 55} (1995), no. 5, 1297--1323.\crr%  %exact solve \De Q_b+Q_b^{p-1}Q_b-ib\Lam Q_b =Q_b 

\bibitem{Kw89} M. Kwong, Uniqueness of positive solutions of $\De u -u + u^p = 0$ in  $\R^n$. {\em Arch. Ration. Mech. Anal}. {\bf 105}  (1989), no.3, 243--266.

%\bibitem{LanPaSuSu88} M.J. Landman, G.C. Papanicolaou, C. Sulem, P.-L. Sulem, Rate of blowup for solutions of the nonlinear Schr\"odinger equation at critical dimension. Phys. Rev. A (3) {\bf 38} (no. 8) (1988) 3837--3843.

%\bibitem{LePaSuSu} B. LeMesurier, G. Papanicolaou, C. Sulem, P. Sulem, Local structure of the self-focusing singularity of the nonlinear Schr\"odinger equation.  Physica D 32 (1988) 210--226.

%\bibitem{LiuZhouZhangZhang} Q. Liu, Y. Zhou, J. Zhang, and W. Zhang, Sharp condition of global existence for nonlinear Schr\"odinger equation with a harmonic potential,  \textit{Appl. Math. Comput.} \textbf{177} (2006), 482--487.


\bibitem{MAHHWC99} M.  Matthews, B.  Anderson, P. Haljan, D. Hall, C. Wiemann, E. A. Cornell, Vortices in a Bose-Einstein condensates. 
{\em Phys. Rev. Lett}. {\bf 83} (1999), 2498--2501.

%\bibitem{Pita61} L.  Pitaevskii, Vortex lines in an imperfect Bose gas. Zh. Eksper. Teor. Fiz. {\bf 40} (1961), 646--651.

\bibitem{Mer} F. Merle,  Determination of blow-up solutions with minimal mass for nonlinear Schr\"odinger equations with critical power.
\textit{Duke Math. J.} \textbf{69} (1993), no. 2, 427--454.

%\bibitem{MeRa02} F. Merle and P. Raphael, Blow up dynamic and upper bound on the blow up rate for critical nonlinear Schr\"odinger equation. In ``Journ\'ees Equations aux D«\'eriv\'ees Partielles'' (Forges-les-Eaux, 2002), pages Exp. No. XII, 5. Univ.  Nantes, Nantes, 2002.

%\bibitem{MeRa03} F. Merle and P. Raphael,  Sharp upper bound on the blow-up rate for the critical nonlinear Schr\"odinger equation. Geom. Funct. Anal., 13(2003):591--642.

\bibitem{MerRa0} F. Merle, P. Rapha\"el,
Blow up dynamic and upper bound on the blow up rate for critical nonlinear Schr\"odinger equation.
\textit{Ann. of Math.} \textbf{161} (2005), no. 1, 157--222.

\bibitem{MerRa1}  F. Merle, P. Rapha\"el,
Profiles and quantization of the blow up mass for critical nonlinear Schr\"odinger equation.
\textit{Comm. Math. Phys.} \textbf{253} (2005), no. 3, 675--704.

%\bibitem{MerRa2} F. Merle and P. Rapha\"el,  On a sharp lower bound on the blow-up rate for the $L^2$-critical nonlinear Schr\"odinger equation,\textit{J. Amer. Math. Soc.} \textbf{19} (2006), no. 1, 37--90.


%\bibitem{MerRa} Frank Merle, Pierre Rapha'l, Igor Rodnianski,  arXiv:1106.0912 
%Blow up dynamics for smooth equivariant solutions to the energy critical Schršdinger map.


%\bibitem{MerRa} Frank Merle, Pierre Raphael, Igor Rodnianski, 
% arXiv:1102.4308     Blow up dynamics for smooth data equivariant solutions to the energy critical Schrodinger map problem.


%\bibitem{MerRaSz} Frank Merle, Pierre Raphael, Jeremie Szeftel,  %arXiv:1010.5168 
%The instability of Bourgain-Wang solutions for the L^2 critical NLS

\bibitem{MeRaSz10} F. Merle, P. Raphael, J. Szeftel, %\href{https://arxiv.org/abs/0907.4098}
 {Stable self similar blow up dynamics for slightly $L^2$ supercritical NLS equations}. {\em Geom. Funct. Anal.}  {\bf 20} (2010), no. 4,  1028--1071.
%(Submitted on 23 Jul 2009)

%We consider the focusing nonlinear Schr\"odinger equations $iu_t+\De u+u|u|^{p-1}=0$ in dimension $1\le N\le 5$ and for slightly L2 supercritical nonlinearities $p_c<p<(1+\eps)p_c$ with $p_c=1+4/N$ and $0<\eps\ll 1$. We prove the existence and stability in the energy space H1 of a self similar finite time blow up dynamics and provide a qualitative description of the singularity formation near the blow up time. 

%	arXiv:0907.4098 [math.AP]

%\bibitem{Mi} L. Michel, Remarks on non-linear Schr\"odinger equation with magnetic fields,  \textit{Comm. Partial Differential Equations} \textbf{33} (2008), no. 7, 1198--1215

%\bibitem{BiSow04} Iwo Bialynicki-Birula, Tomasz Sowinski (Center for Theoretical Physics PAS),\href{https://arxiv.org/abs/quant-ph/0310195}{Solutions of the Logarithmic Schrodinger Equation in a Rotating Harmonic Trap.}(Submitted on 31 Oct 2003 (v1), last revised 13 Sep 2004 (this version, v3))

%We study the influence of the nonlinearity in the Schrodinger equation on the motion of quantum particles in a harmonic trap. In order to obtain exact analytic solutions, we have chosen the logarithmic nonlinearity. The unexpected result of our study is the existence in the presence of nonlinearity of two or even three coexisting Gaussian solutions.
%	8 pages, 3 figures	Quantum Physics (quant-ph)
	%arXiv:quant-ph/0310195 	%(or arXiv:quant-ph/0310195v3 for this version)


\bibitem{Pere01} G. Perelman, On the blow up phenomenon for the critical nonlinear 
Schr\"odinger equation in 1D. {\em Ann. Henri Poincar\'e} {\bf 2} (2001), 605--673. 

% P. Raphael, Stability of the log-log bound for blow up solutions to the critical nonlinear Schrodinger equation, Math. Ann. 331 (3) (2005), 577--609
\bibitem{Raph05} P. Raphael, %\href{https://link.springer.com/article/10.1007/s00208-004-0596-0}
{Stability of the $\log$-$\log$ bound for blow up solutions to the critical nonlinear Schr\"odinger equation.} {\em Math. Ann.}  {\bf  331} (2005), no. 3, 577--609.
%First Online: 12 January 2005.%\edz{Jan.22 Fan said that the stability in $H^1$ should be due to Raphael 05?} 

%Abstract.  We consider finite time blow up solutions to the critical nonlinear Schrodinger equation Open image in new window with initial condition $u_0\in H^1$. Existence of such solutions is known, but the complete blow up dynamic is not understood so far. For initial data with negative energy, finite time blow up with a universal sharp upper bound on the blow up rate corresponding to the so-called log-log law has been proved in [10], [11]. We focus in this paper onto the positive energy case where at least two blow up speeds are known to possibly occur. We establish the stability in energy space $H^1$ of the log-log upper bound exhibited in the negative energy case, and a sharp lower bound on blow up rate in the other regime which corresponds to known explicit blow up solutions.


\bibitem{ReZaStri01} A. Recati, F. Zambelli,  S. Stringari,  Overcritical rotation of a trapped Bose-Einstein condensate. 
{\em Phys. Rev. Lett}. {\bf 86} (2001), 377--380.  %Published 15 January 2001 

%\iffalse The rotational motion of an interacting Bose-Einstein condensate confined by a harmonic trap is investigated by solving the hydrodynamic equations of superfluids, with the irrotationality constraint for the velocity field. We point out the occurrence of an overcritical branch where the system can rotate with angular velocity larger than the oscillator frequencies. We show that in the case of isotropic trapping the system exhibits a bifurcation from an axisymmetric to a triaxial configuration, as a consequence of the interatomic forces. The dynamical stability of the rotational motion with respect to the dipole and quadrupole oscillations is explicitly discussed.

%[from the article and sz]: Give a dilute Bose gas interacting with repulsive forces at zero temperature, for large systems the equations of motion are  described by the hydrodynamical theory of superfluids. When the anisotropic harmonic trapping potential rotates with angular velocity $\Om$, these equation takes the form in the corotating frame:
%\begin{align}
%& \frac{\pa \rho}{\pa t}+\nabla (\rho ({\bf u}-\Om \times {\bf r}))=0\label{e:rho-u-Om}\\
%& \frac{\pa u}{\pa t}+\nabla \cdot ( \frac{\nu^2}{2}+ \frac{V}{M} +\frac{\mu}{M}- {\bf u}\cdot (\Om \times {\bf r}))=0\label{e:ut-VOm}
%\end{align}
%where $V=V_{ext}({\bf r})=M(\ga_x^2 x^2+\ga_y^2y^2+\ga_z^2z^2)/2$ is the (time-independent) oscillator potential providing the external confinement (with $\ga_x>\ga_y$).
%$\mu=\mu_{loc}(\rho)=g\rho$ is the chemical potential of the uniform gas, here $g=4\pi \hbar^2 a/M$ is the coupling constant fixed by the $s$-wave scattering length $a$ and 
%${\bf u}$ is the velocity field %in the laboratory frame,  expressed in terms of the coordinates in the rotating frame. It satisfies the {\em irrotationality constraint}.  Equations \ref{e:rho-u-Om}, \ref{e:ut-VOm} also apply to a trapping Fermi superfluid where $\mu$ takes a different form. \crr Vortices cannot be in any case described by the hydrodynamical equations \ref{e:rho-u-Om}, \ref{e:ut-VOm} since they require the use of more microscopic approaches. GPE \eqref{eq:nls-u} can provide such a model. %theory.  

%(since the critical frequency needed to generate a stable vortex becomes smaller and smaller as the number of atoms increases, the solutions in \cite{ReZaStri01} corresponds to metastable configurations in general.)
%\fi

 %K. W. Madison et al., Phys. Rev. Lett. 86, 4443 (2001).

% S. Sinha and Y. Castin, Phys. Rev. Lett. 87, 190402  (2001).

% J. J. Garcia-Ripoll and V. M. Perez-Garcia, Phys. Rev. A 64, 013602 (2001).

%\bibitem{Ros02} P. Rosenbusch et al., Phys. Rev. Lett. 88, 250403 (2002)

%J. Colliander and P. Raphael. Rough blowup solutions to the L2 critical NLS.  Math. Ann., 345(2009):307-366.


%\bibitem{SowBi04} Tomasz Sowinski, Iwo Bialynicki-Birula, %(Center for Theoretical Physics PAS), 
 %\href{https://arxiv.org/abs/quant-ph/0409070}
 %{Harmonic oscillator in a rotating trap: Complete solution in 3D}.  (2004).

%Complete description of the classical and quantum dynamics of a particle in an anisotropic, rotating, harmonic trap is given. The problem is studied in three dimensions and no restrictions on the geometry are imposed. In the generic case, for an arbitrary orientation of the rotation axis, there are two regions of instability with different characteristics. The analysis of the quantum-mechanical problem is made simple due to a direct connection between the classical mode vectors and the quantum-mechanical wave functions. This connection is obtained via the matrix Riccati equation that governs the time evolution of squeezed states of the harmonic oscillator. It is also shown that the inclusion of gravity leads to a resonant behavior -- the particles are expelled from the trap.
%Comments:	6 figures	Quantum Physics (quant-ph)	arXiv:quant-ph/0409070
% 	(or arXiv:quant-ph/0409070v1 for this version)


%\bibitem{Tao09} Tao, Terence,  \href{http://emis.ams.org/journals/NYJM/j/2009/15-14.pdf}{A pseudoconformal compactification of the nonlinear Schr\"odinger equation and applications.}\crr   New York J. Math 15 (2009): 265--282.

\bibitem{Wein83} M. Weinstein, Nonlinear Schr\"odinger equations and sharp interpolation estimates. {\em Comm. Math. Phys}. {\bf 87} (1983) 567--576. 

\bibitem{Ya} K. Yajima, 
Schr\"odinger evolution equations with magnetic fields.
\textit{J. Analyse Math.} \textbf{56} (1991), 29--76. 

\bibitem{Zheng}
S. Zheng, Fractional regularity for nonlinear Schr\"odinger equations with magnetic fields.
\textit{Contemp. Math.} \textbf{581} (2012), 271--285.

\bibitem{ZhuZhang}
S. Zhu,  J. Zhang, Profiles of blow-up solutions for the Gross-Pitaevskii equation.
\textit{Acta Math. Appl. Sin.} \textbf{26} (2010), no. 4, 597--606.
\end{thebibliography}

\documentclass[12pt, reqno]{amsart}
\usepackage{amscd}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage{amsxtra,hyperref}
\usepackage{array}
\usepackage{enumitem}
\usepackage[dvips]{graphicx}
\usepackage{latexsym}

\usepackage[mathscr]{eucal}
\usepackage{eucal} 

\usepackage{graphicx,psfrag,color}
\usepackage{epsfig,pstricks}
 \usepackage{mathrsfs}
\usepackage{calligra}
\usepackage[T1]{fontenc}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}

\theoremstyle{definition}
\newtheorem{remark}[theorem]{Remark}

\allowdisplaybreaks

\newtheorem*{SP}{Spectral~Property} 

\definecolor{orange}{rgb}{0.995, 0.75, 0.35}
\definecolor{purple}{rgb}{0.7, 0.2, 0.5}
\definecolor{royalblue}{rgb}{0.2, 0.7, 0.8}
\definecolor{darkgreen}{rgb}{0.2,0.725,0.25} 

\newcommand{\norm}[1]{ \left|  #1 \right|}

\def\al{\alpha}
\def\de{\delta}

\def\ga{\gamma}
\def\veps{\varepsilon}
\def\vphi{\varphi}

\def\De{\Delta}

\def\Lam{\Lambda}
\def\Om{\Omega}

\def\dive{\mathrm{div}\,}

\def\iy{\infty}

\def\pa{\partial}

\def\sH{\mathscr{H}}

\def\trace{\mathrm{tr}\;}

\newcommand{\inv}{^{-1}}

\newcommand{\nd}{\noindent}

\newcommand{\R}{\mathbb{R}}
\newcommand{\C}{\mathbb{C}}

\def\sideremark#1{\ifvmode\leavevmode\fi\vadjust{\vbox to0pt{\vss
 \hbox to 0pt{\hskip\hsize\hskip1em
\vbox{\hsize2cm\tiny\raggedright\pretolerance10000
 \noindent #1\hfill}\hss}\vbox to8pt{\vfil}\vss}}}%

{\end{list}}

\begin{document}
\title[Blowup rate for rotational NLS]{Blowup rate for mass critical rotational nonlinear Schr\"odinger equations}
\author{Nyla Basharat}
\address[Nyla Basharat]{
DPHS\\
Augusta University\\
Augusta, GA 30912}
\email{nbasharat@augusta.edu}
\author{Yi Hu}
\address[Yi Hu]{Department of Mathematical Sciences \\
Georgia Southern University \\
Statesboro, GA 30460}
\email{yihu@georgiasouthern.edu}
\author{Shijun Zheng}
\address[Shijun Zheng]{DMS\\
Georgia Southern University \\
Statesboro, GA 30460}
\email{szheng@georgiasouthern.edu}

\date{\today}
\subjclass[2010]{Primary  35Q55, 35B44. Secondary 35P30}
\keywords{blowup rate, harmonic potential, angular momentum}

\begin{abstract}
We consider the blowup rate for blowup  solutions to $L^2$-critical, focusing NLS 
with a harmonic potential and a rotation term.
Under a suitable spectral condition we  prove that there holds the ``$\log$-$\log$ law'' when the initial data is slightly above the ground state. 
We also construct  minimal mass blowup solutions near the ground state level with distinct blowup rates. 
\end{abstract}

\maketitle


\section{Introduction}
 Consider the focusing nonlinear Schr\"odinger equation (NLS) with an angular momentum term on $\mathbb{R}^{1+n}$:
	\begin{align}\label{eq:nls-u}
	\begin{cases}
	i u_t = - \Delta u + V u - |u|^{p-1} u + i A \cdot \nabla u \\
	u(0, x)=u_0\in \sH^1,
	\end{cases}
	\end{align}
where $u = u(t, x)$: $\mathbb{R} \times \mathbb{R}^n\to\C$ denotes the wave function, 
$V(x)= \gamma^2 |x|^2$, $\gamma>0$ is a trapping harmonic potential that confines the movement of particles,  
and $A(x) = M x$, 
with $M=-M^T$ being an $n \times n$ real-valued skew-symmetric matrix.  
The linear hamiltonian $H_{A, V} := -\Delta + V + i A \cdot \nabla$ is essentially self-adjoint in $L^2$,
whose eigenvalues are associated to the Landau levels as quantum numbers. 
The angular momentum operator $L_A u:=i A\cdot\nabla u$  generates the rotation $e^{tA\cdot \nabla }u=u(e^{tM}x)$ in $\mathbb{R}^n$. 
The space $\sH^1=\sH^{1,2}$ denotes the weighted Sobolev space given by: for $r\in (1,\iy)$,
	\begin{align*}	\mathscr{H}^{1,r}(\mathbb{R}^n)
	:= \left\{ f \in L^r:  \nabla f,\,  x f \in L^r\right\},
	\end{align*}
which is endowed with the norm $\norm{f}_{\sH^{1,r}}=\norm{\nabla f}_r+\norm{x f}_r+\norm{ f}_r$, here  
$\norm{\cdot}_r:=\norm{\cdot}_{L^r}$ denoting the $L^r$-norm. 

When $n =3$, equation \eqref{eq:nls-u} is also known as  Gross-Pitaevskii equation which models rotating Bose-Einstein condensation (BEC) with attractive particle interactions in a dilute gaseous ultra-cold superfluid.  
The operator $L_A$ are usually denoted by $- \Omega \cdot L $,
where $\Omega=(\Om_1,\Om_2,\Om_3) \in \mathbb{R}^3$ is a given angular velocity vector and $L := -i x \wedge \nabla$, in which case
$M= \begin{pmatrix} 0& -\Om_3& \Om_2\\
\Om_3&0& -\Om_1\\   
-\Om_2&\Om_1&0
\end{pmatrix}$.  

Such a system describing rotating particles in a 
harmonic trap has acquired significance in connection with optics and atomic physics 
in theoretical and experimental physics \cite{Gross61,MAHHWC99,ReZaStri01,
BaoWMar05,Af06,BaoCai15}. 
Meanwhile, it demands rigorous mathematical analysis on the evolution and dynamics of the quantized flow. 
For $p\in (1, 1+4/(n-2))$, the local in time existence and uniqueness of (\ref{eq:nls-u}) has been established in  \cite{HaoHsiaoLi1,HaoHsiaoLi2,AnMaSpar,BaHaHuZheng},
see also \cite{CazE88,De,Zheng} for the treatment in a general magnetic setting.  
The purpose of this article is to study how the rotation  affects the  wave collapse as well as energy concentration under a  trapping potential.
In particular, we will address the blowup rate for the blowup solution of the $L^2$-critical focusing equation in \eqref{eq:nls-u} where $p=1+4/n$. 

Let $H^1=\{u\in L^2:\nabla u\in L^2\}$ be the usual Sobolev space. Recall that the standard NLS reads 
	\begin{align}\label{eq:nls-phi}
	\begin{cases}
	i \varphi_t = - \Delta \varphi - |\varphi|^{p-1} \varphi, \\
	\varphi(0, x) = \varphi_0\in H^1,
	\end{cases}
	\end{align}
and the well-posedness problem for $p\in (1,1+4/(n-2)]$ has been studied for a few decades and is quite well understood in the euclidean space.
Let $Q \in H^1(\mathbb{R}^n)$ be the unique   positive radial function that satisfies (\cite{Kw89,Wein83})
	\begin{align}\label{eq:ground-state}
	-Q = -\Delta Q - Q^{1+\frac{4}{n}}.
	\end{align}
When $p=1 + \frac{4}{n}$  and $\norm{\vphi_0}_2=\norm{Q}_2$,
Merle \cite{Mer} was able to determine the profile for all blowup solutions with minimal mass at the ground state level, which are obtained from pseudo-conformal transform. 
Hence
 all blowup solutions have blowup rate $(T-t)\inv$. 
In the mass critical and supercritical case $p\in [1+4/n,\iy)$, the wave collapse dynamics appears very subtle issue for $\norm{u_0}_2=\norm{Q}_2$.  
Within an arbitrarily small neighborhood of $Q$, there always exist $\vphi_0$ and $\psi_0$ such that the flow 
$\vphi_0\mapsto \vphi$ blowups in finite time, and, $\psi_0\mapsto \psi$ exists global in time and scatters as $t\to \iy$ in $H^1$. 

When $p=1+\frac{4}{n}$ and $\norm{\vphi_0}_2>\norm{Q}_2$, Bourgain and Wang \cite{BW98} 
constructed solutions of positive energy having blowup rate $(T-t)\inv$ in dimensions $n=1, 2$, 
which  was later shown unstable however.  
Perelman \cite{Pere01}  gave the first rigorous demonstration of the existence and stability for the $\log$-$\log$ speed for generic blowup solutions in 1d. 
More recently, under certain spectral condition in the Spectral Property (Section \ref{s:spec-loglog}), 
Merle and Rapha\"el \cite{MerRa1} 
proved the sharp blowup rate of the solutions for  \eqref{eq:nls-phi},
i.e., there exists a universal constant $\al^*$ such that 
if $\varphi_0\in B_{\al^*}$ with negative energy,
then  $\vphi(t,x)$ is a blowup solution to \eqref{eq:nls-phi} with maximal interval of existence $[0, T)$
 satisfying the $\log$-$\log$ blowup rate as $t\to T$
	\begin{align*}
         \vert \nabla \varphi(t) \vert_{2}
	\approx\sqrt{ \frac{\log |\log(T - t)|}{T-t} },
	\end{align*}
where
	\begin{align*}
	  B_{\alpha}
	:= \left\{ \phi \in H^1: \int |Q|^2 dx < \int |\phi|^2 dx < \int |Q|^2 dx + \alpha \right\},
	\end{align*}
see Theorem \ref{thm:log-log-phi}.
Such a blowup rate is also known to be stable in $H^1$.  

In the presence of a rotational term,  we will show how to prove   such a ``$\log$-$\log$ law''  
for the NLS  \eqref{eq:nls-u}.   
Our proof is based on a virial identity for \eqref{eq:nls-u}, 
 the $\mathcal{R}$-transform \eqref{eq:phi-to-u} that maps solutions of \eqref{eq:nls-phi} to solutions of \eqref{eq:nls-u}, and an application of the above result of Merle and Rapha\"el's. 
 This treatment is motivated by  \cite{ZhuZhang}, where the analogous result is obtained for the case $A=0$ and $V$ being a harmonic potential.  
Our main result is stated as follows. Let $\al^*$ be the above-mentioned universal constant. 
\begin{theorem}\label{thm:log-log-u}
Let $p = 1 + \frac{4}{n}$ in \eqref{eq:nls-u},  $1\le n \leq 5$. 
Suppose $u_0\in B_{\al^*}$ satifies
	\begin{align}\label{eq:negative-initial-energy}
	\int |\nabla u_0|^2 -\frac{n}{n+2} \int |u_0|^{2+\frac{4}{n}}
	< 0.
	\end{align}
Then $u\in C([0, T); \mathscr{H}^1)$ is a blowup solution of  \eqref{eq:nls-u} in finite time $T < \infty$, 
with the $\log$-$\log$ blowup rate
	\begin{align*}
	\lim_{t \rightarrow T}
	\frac{\vert \nabla u(t) \vert_2}{\vert \nabla Q \vert_{2}}
	\sqrt{ \frac{T - t}{\log \left| \log (T - t)  \right|} }
	= \frac{1}{\sqrt{2 \pi}},
	\end{align*}
where $Q$ is the unique solution of \eqref{eq:ground-state}.
\end{theorem}

\section{Local wellposedness for \eqref{eq:nls-u}} 
For $\varphi_0 \in \mathscr{H}^1$, 
the local well-posedness of \eqref{eq:nls-u} was proved for $1 \leq p < 1 + \frac{4}{n-2}$,  
see e.g. in \cite{De, Zheng}.  %
  The proof for the local well-posedness relies  on local in time dispersive estimates for $U(t)=e^{-itH_{A,V}}$,  
the fundamental solution on $[0,\de)$ (for some small $\de>0$) constructed in \cite{Ya}. 
The vectorial function $A$  represents a magnetic potential that induces the Coriolis effect or centrifugal force for the spinor particles. 
Alternatively, 
this can also be done by means of the  explicit formula in \eqref{e:fund-U(t)} for $U(t)$ defined in $(0,\pi/2\ga)$. 
This formula is obtained from the so-called $\mathcal{R}$-transform in \eqref{eq:phi-to-u}, a type of pseudo-conformal transform.

\begin{proposition}\label{p:blup-altern-conserv} 
Let $1 + \frac{4}{n} \leq p < 1 + \frac{4}{n-2}$. Suppose $u_0 \in \mathscr{H}^1$.  
\begin{enumerate}
\item[\textup{(a)}]\;\textup{[Blowup Alternative]}  
	\begin{enumerate}[label=\rm{(\alph*)}]
	\item[\textup{(i)}] If $1 \leq p < 1 + \frac{4}{n-2}$,  then there exists $T^* > 0$ such that \eqref{eq:nls-u} has a unique maximal solution $u \in C([0, T^*), \mathscr{H}^1)\cap L^q_{loc}([0, T^*), \mathscr{H}^{1,r})$, where $r = p+1$  and $q= \frac{4(p+1)}{n(p-1)}$.
\item[\textup{(ii)}] If $T^*$ is finite, then $\norm{\nabla u}_2\to \iy$  as $t\to T^*$ with a lower bound:
\begin{align*}
&\norm{\nabla u(t)}_2\ge \frac{C}{\sqrt{T^*-t}}.
\end{align*}
\end{enumerate}	
	\item[\textup{(b)}]\;\textup{[Conservation Laws]} The following are conserved on the maximal lifespan $[0,T^*)$.
		\begin{enumerate}[label=\rm{(\roman*)}]
		\item[]\;\textup{(mass)}
		$\displaystyle M(u)=\int |u|^2$
	\item[]\;\textup{(energy)}
		$\displaystyle E(u)=\int \left( |\nabla u|^2 +V |u|^2 - \frac{2}{p+1} |u|^{p+1} + i \bar{u} A\cdot \nabla u \right)$
	\item[]\;\textup{(angular momentum)}
		$\displaystyle \ell_A(u)=\int i \bar{u} A\cdot \nabla u$.
		\end{enumerate}
\end{enumerate} 
\end{proposition} 
In the critical case $p=1+4/n$, from \cite{BaHaHuZheng} we know that  $\norm{Q}_2$ is the sharp threshold such that: 
\begin{enumerate}
\item[(a)] If $\norm{u_0}<\norm{Q}_2$, then \eqref{eq:nls-u} has a unique global in time solution.
\item[(b)] For all $c\ge \norm{Q}_2$, there exists $u_0$ with   $\norm{u_0}_2=c$ so that $u$ is a finite time blowup solution of \eqref{eq:nls-u}.
\end{enumerate}
As we mentioned in the introduction section, if $\norm{u_0}=\norm{Q}_2$, 
 from Merle's characterization for the blowup profile of \eqref{eq:nls-phi}, all such blowup solutions have the blowup rate 
 $\norm{\nabla u(t)}_2\approx \frac{C}{T^*-t}$ as $t\to T^*$, see Proposition  \ref{c:charact-mini-pc}.

\section{A spectral property and the $\log$-$\log$ law}\label{s:spec-loglog} 
To  show the blowup rate for initial data above the ground state $Q$ 
as stated in  Theorem \ref{thm:log-log-u},
 we need the following Spectral Property.
Let  $y$ denote the spatial variable in $\R^n$.
\begin{SP}\label{h:spectral_property}
Consider the two  Schr\"odinger operators
	\begin{align*}
	L_1
	:= - \Delta + \frac{2}{n} \left( \frac{4}{n} + 1 \right) Q^{\frac{4}{n} - 1} y \cdot \nabla Q, \qquad
	L_2
	:= - \Delta + \frac{2}{n} Q^{\frac{4}{n} - 1} y \cdot \nabla Q,
	\end{align*}
and the real-valued quadratic form for $\varepsilon = \varepsilon_1 + i \varepsilon_2 \in H^1$ 
	\begin{align*}
	H(\varepsilon, \varepsilon)
	:= (L_1 \varepsilon_1, \varepsilon_1) + (L_2 \varepsilon_2, \varepsilon_2).
	\end{align*}
Let
	\begin{align*}
	Q_1
	:= \frac{n}{2} Q + y \cdot \nabla Q, \qquad
	Q_2
	:= \frac{n}{2} Q_1 + y \cdot \nabla Q_1.
	\end{align*}
Then there exists a universal constant $\delta_0 > 0$ such that for every $\varepsilon \in H^1$,
if
	\begin{align*}
	(\varepsilon_1, Q)
	= (\varepsilon_1, Q_1)
	= (\varepsilon_1, y_j Q)_{1 \leq j \leq n}
	= (\varepsilon_2, Q_1)
	= (\varepsilon_2, Q_2)
	= (\varepsilon_2, \partial_{y_j} Q)_{1 \leq j \leq n}
	= 0,
	\end{align*}
then
	\begin{align*}
	H(\varepsilon, \varepsilon)
	\geq \delta_0 \left( \int |\nabla \varepsilon|^2 dy + \int |\varepsilon|^2 e^{-|y|} dy \right).
	\end{align*}
\end{SP}

The proof of the Spectral Property in any dimension is not complete.
It has been proved in \cite{MerRa0} for dimension $n=1$ by using the explicit solution
$Q(x) = \left( \frac{3}{\cosh^2(2x)} \right)^\frac{1}{4}$ to  \eqref{eq:ground-state}. 
One can also find a computer assisted proof of the Spectral Property in dimensions $n = 2, 3, 4$ in \cite{FiMerRa}.
For dimension $5$ and higher see Remark 
\ref{re:N56_orthogonal}.

 The  Spectral Property is equivalent to the coercivity for $L_1$ and $L_2$ on quadratic forms,
 the study of which involving the ground
state solution $Q$ naturally appears in a perturbation setting when dealing with stability problem. 
These two operators are related to the Lyapounov functionals $L_\pm$,  where 
$L_+=-\De+1 -(1+\frac{4}{n})  Q^{\frac{4}{n}} $ and
$L_-=-\De+1 -Q^{\frac{4}{n}} $, see \cite{FiMerRa}. 


Using the Spectral Property, 
Merle and Rapha\"el obtained the following blowup rate for  \eqref{eq:nls-phi} in the absence of potentials,
see  \cite{MerRa1,FiMerRa,MeRaSz10}. 
\begin{theorem}\label{thm:log-log-phi}
Let $p = 1 + \frac{4}{n}$ in \eqref{eq:nls-phi}. 
Let $1\le n \leq 5$. 
 There exists a universal constant $\alpha^* > 0$ such that the following is true.
 Suppose	$\varphi_0 \in B_{\alpha^*}$ 
satisfies
	\begin{align*}
	\int |\nabla \varphi_0|^2 -\frac{n}{n+2} \int |\varphi_0|^{2+\frac{4}{n}}
	< 0.
	\end{align*}
Then  $\varphi \in C([0, T); H^1)$ is a blowup solution of  \eqref{eq:nls-phi} in finite time $T < \infty$,
which  admits the $\log$-$\log$ blowup rate
	\begin{align}\label{e:log-log_Q}
	\lim_{t \rightarrow T}
	\frac{\vert \nabla \varphi(t) \vert_2}{\vert \nabla Q \vert_2}
	\sqrt{ \frac{T - t}{\log \left| \log (T - t)  \right|} }
	= \frac{1}{\sqrt{2 \pi}}.
	\end{align}
\end{theorem}

\begin{remark}\label{re:N56_orthogonal} 
From \cite{FiMerRa} we know that the Spectral Property is true in dimensions $n=1,2,3,4$.
 If  $n= 5$, Theorem \ref{thm:log-log-phi} continues to hold true as soon as
  the Spectral Property verifies 
  the orthogonality condition 
  with  $(\veps_1, Q_1) = 0$ replaced by $(\veps_1, |y|^2Q) = 0$, which is numerically verified in \cite{FiMerRa}. 
For the above-mentioned reason, Theorem \ref{thm:log-log-phi} remains open in dimensions $n\ge 6$.
\end{remark}

\begin{remark} It is well-known that the $\log$-$\log$ law is {a} generic behavior for those blowup solutions  in the theorem,
whose proof relies on  algebraic cancellations  related to the topological degeneracy of the linear 
operators $L_1$ and  $L_2$ around $Q$. 
Such blowup rate is stable in the sense that
the set $\mathcal{U}_0$ in the log-log regime is open in $H^1$, where 
$\mathcal{U}_0$ denotes the set of all initial
data $\vphi_0$ in $B_{\al^*}$ 
 so that the flow  $\vphi_0\mapsto \vphi(t)$ of (\ref{eq:nls-phi})
collapses  in finite time $T^* <\iy$ with the $\log$-$\log$ speed given in (\ref{e:log-log_Q}), see \cite{Raph05,FiMerRa}.
\end{remark}

\section{The blowup rate 
for the rotational NLS}
In this section we  prove Theorem \ref{thm:log-log-u}.
We will always assume $p = 1 + \frac{4}{n}$ in both \eqref{eq:nls-u} and \eqref{eq:nls-phi}. 
We will need a  virial identity for \eqref{eq:nls-phi} and the $\mathcal{R}$-transform introduced in Proposition \ref{p:u-phi-R_pseudo-conformal}. 
This  transform  gives a relation between the two solutions of  \eqref{eq:nls-u} and \eqref{eq:nls-phi},
which is  coined as a combination of the lens transform and the rotation $e^{t A\cdot \nabla}$. One  can view it as certain pseudo-conformal symmetry in the rotational case, see \cite{Car,ZhuZhang} in the presence of a quadratic potential only, i.e., $M=0$ and $\ga\ne 0$. 

The following is a standard virial identity for \eqref{eq:nls-phi} in the weighted Sobolev space $\sH^1$, which can be proved by a direct calculation. 
\begin{lemma}\label{lem:virial-identity} 
Let $\varphi$ be a solution to the problem \eqref{eq:nls-phi} in $C([0, T), \sH^1)$. 
Define $J(t) := \int |x|^2 |\varphi|^2 dx$. Then
	\begin{align*}
	J'(t)
	= 4 \Im \int x \overline{\varphi} \cdot \nabla \varphi dx, \qquad
	J''(t)
	= 8 \mathcal{E}(\varphi_0),
	\end{align*}
where
	\begin{align*}
	\mathcal{E}(\varphi) = \int \left( |\nabla \varphi|^2 - \frac{n}{n+2} |\varphi|^{\frac{4}{n} + 2} \right) dx.
	\end{align*}
\end{lemma}



\begin{lemma}\label{lem:e^M}
\begin{enumerate}
\item[\textup{(a)}] Given any real $n$ by $n$ matrix $M$  and any  function $f$ in $C_0^\iy\cap L^2(\R^n)$, 
we have
	\begin{align}\label{E:etMx}
	e^{t (M x) \cdot \nabla} f(x)
	= f(e^{tM} x).
	\end{align}
\item[\textup{(b)}] If $M$ is a real skew-symmetric matrix, then $e^{tM}\in SO(n)$ for all $t$, where $SO(n)$ is the group of $n$ by $n$ orthogonal matrices with determinant $1$. 
Moreover, $(Mx) \cdot \nabla$ and $\Delta$ commute,
i.e.,
	\begin{align*}
	[(Mx) \cdot \nabla, \Delta] = 0.
	\end{align*}
\end{enumerate}\end{lemma} 
Part (a) can be proven by showing that both sides of \eqref{E:etMx} obey the ODE
\begin{align*}
	\partial_t F(t, x)= (Mx) \cdot \nabla F(t, x),  \quad F(0, x) = f(x).
	\end{align*}
Part (b) follows from a straightforward calculation.

\begin{proposition}\label{p:u-phi-R_pseudo-conformal}
Let $\varphi(t,x) \in C([0,T), H^1)$ be a solution to \eqref{eq:nls-phi} where $T > 0$.
Define the $\mathcal{R}$ transform $\vphi\mapsto \mathcal{R}(\vphi)$ to be
	\begin{align}\label{eq:phi-to-u}
	\mathcal{R}\vphi(t, x)
	:= \frac{1}{\cos^\frac{n}{2}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	\varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right).
	\end{align}
Then $u=\mathcal{R}\vphi$ is a solution to \eqref{eq:nls-u} in $C([0, \frac{\arctan(2 \gamma T)}{2 \gamma}), \mathscr{H}^1)$.

Conversely, let $u(t,x) \in C([0,T^*), \mathscr{H}^1)$
be a solution to \eqref{eq:nls-u} where $T^* \in (0, \frac{\pi}{4 \gamma}]$.
Then $\vphi=\mathcal{R}\inv u$,  given by 
	\begin{align}\label{eq:u-to-phi}
	\varphi(t, x)
	:= \frac{1}{(1 + (2 \gamma t)^2)^\frac{n}{4}}
	e^{i \frac{\gamma^2 |x|^2 t}{1 + (2 \gamma t)^2}}
	u\left( \frac{\arctan(2 \gamma t)}{2 \gamma}, \frac{e^{-tM} x}{\sqrt{ 1 + (2 \gamma t)^2}} \right),
	\end{align}
is a solution to \eqref{eq:nls-phi} in $C([0, \frac{\tan(2 \gamma T^*)}{2 \gamma}), H^1)$, where $\mathcal{R}\inv $ is the inverse of $\mathcal{R}$.
\end{proposition}

\begin{proof} We will only briefly check \eqref{eq:phi-to-u} for $u=\mathcal{R}\phi$.
The other one $u\mapsto \vphi$ is the inverse transform. 
By direct computation,
we have
	\begin{align}\label{eq:u_t}
	\begin{split}
	u_t(t,x)
	& = n \gamma \frac{\sin(2 \gamma t)}{\cos^{\frac{n}{2}+1}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	\varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \\
	& \quad - i \gamma^2 |x|^2 \frac{1}{\cos^{\frac{n}{2}+2}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	\varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \\
	& \quad + \frac{1}{\cos^{\frac{n}{2}+2}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	\varphi_t\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \\
	& \quad + \frac{1}{\cos^{\frac{n}{2}+1}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	(e^{tM} M x) \cdot \nabla \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \\
	& \quad + 2 \gamma \frac{\sin(2 \gamma t)}{\cos^{\frac{n}{2}+2}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	(e^{tM} x) \cdot \nabla \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right).
	\end{split}
	\end{align}
To compute $\Delta u$ and $i (Mx) \cdot \nabla u$,
first note that
	\begin{align*}
	& \nabla\left( \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \right)
	= \frac{1}{\cos(2 \gamma t)} (e^{tM})^T \nabla \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right), \\
	& \Delta\left( \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \right)
	= \frac{1}{\cos^2(2 \gamma t)} \Delta \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right),
	\end{align*}
where we used 
\begin{enumerate} 
\item[(a)] If $B\in M_{n\times n}$ is a constant matrix, then 
$\nabla (\vphi(B x) )=B^T (\nabla \vphi)(Bx)$;
\item[(b)] $\De=\dive(\nabla)$, $\dive{{\bf F}}=\mathrm{tr}(\frac{\pa (F_1\cdots,F_n)}{\pa(x_1,\cdots,x_n)})=\text{trace of the Jacobian of}\;\mathbf{F} $;
\item[(c)] If $C$ is a constant square matrix, $\mathbf{W}=[w_1,\cdots,w_n]^T$ is a vector-valued function of $x\in\R^n$, then
\begin{align*}
&\dive(C \mathbf{W})=\trace([ \nabla w_1, \cdots, \nabla w_n] C^T).
\end{align*}
\item[(d)] $\trace(U^* \Lam U)=\trace(\Lam)$ if $U$ is a unitary matrix and $\Lam\in M_{n\times n}$.
\end{enumerate}
Thus  we obtain
	\begin{align}\label{eq:laplacian-u}
	\Delta u(t,x)
	& = -i n \gamma \frac{\sin(2 \gamma t)}{\cos^{\frac{n}{2}+1}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	\varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \notag\\
	& \quad - \gamma^2 |x|^2 \frac{\sin^2(2 \gamma t)}{\cos^{\frac{n}{2}+2}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	\varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \notag\\
	& \quad -i 2 \gamma \frac{\sin(2 \gamma t)}{\cos^{\frac{n}{2}+2}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	(e^{tM} x) \cdot \nabla\varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \notag\\
	& \quad + \frac{1}{\cos^{\frac{n}{2}+2}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	\Delta \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right),
	\end{align}
and, noting that $(Mx) \cdot x = 0$, 
	\begin{align}\label{eq:imxdu}
	i (Mx) \cdot \nabla u
	 =  \frac{i}{\cos^{\frac{n}{2}+1}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	(M e^{tM} x) \cdot \nabla  \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right).
	\end{align}
Bring \eqref{eq:phi-to-u},
\eqref{eq:u_t},
\eqref{eq:laplacian-u},
and \eqref{eq:imxdu} into Cauchy problem \eqref{eq:nls-u},
and recall that $\varphi$ satisfies \eqref{eq:nls-phi} with $p=1+4/n$,
hence, we see that $u$ is a solution to \eqref{eq:nls-u} with $u(0,x)=\vphi(0,x)$. 

The above virtually shows that under the relation $u=\mathcal{R}\vphi\iff \vphi=\mathcal{R}\inv u$,
$u$ satisfies  \eqref{eq:nls-u} if and only if $\vphi$ satisfies  \eqref{eq:nls-phi}.  Therefore the second part of the proposition is also true.
\end{proof}
\begin{remark} The $\mathcal{R}$ transform also allows us to solve the linear equation for \eqref{eq:nls-u}.  
The equation 
$i\pa_t\vphi=-\De \vphi $ has the fundamental solution
	\begin{align*}
	e^{i t \Delta} (x,y)
	= \frac{1}{(4 \pi i t)^\frac{n}{2}} e^{i \frac{|x - y|^2}{4t}} .
	\end{align*}
Applying \eqref{eq:phi-to-u}  we then  obtain the fundamental solution to $i\pa_t u=H_{A, V}u$: 
	\begin{align}\label{e:fund-U(t)}
	 e^{-itH_{A,V}}(x,y)=\left( \frac{\gamma}{2 \pi i \sin(2 \gamma t)} \right)^\frac{n}{2}
	e^{i \frac{\gamma}{2} (|x|^2+|y|^2) \cot(2 \gamma t)}
	e^{-i \gamma \frac{(e^{tM} x) \cdot y}{\sin(2 \gamma t)}}.	
	\end{align}
This expression 
 is significantly simpler than the one  in \cite{ChMarS16} per Mehler's formula. 
  To our best knowledge,  \eqref{e:fund-U(t)} might be  the first simply unified explicit formula compared with 
  \cite{Kita80,HaoHsiaoLi1, HaoHsiaoLi2, AnMaSpar}. 
\end{remark}

Now we are ready to prove Theorem \ref{thm:log-log-u}.
\begin{proof}[Proof of Theorem \ref{thm:log-log-u}]
Let $u \in C([0, T), \mathscr{H}^1)$ be the blowup solution to the  problem \eqref{eq:nls-u},
where $[0, T)$ is the maximal interval of existence.
Then by \eqref{eq:u-to-phi},
there is a $\varphi(t,x) \in C([0, \frac{\tan(2 \gamma T)}{2 \gamma}), \mathscr{H}^1)$ that solves  \eqref{eq:nls-phi},
where $[0, \frac{\tan(2 \gamma T)}{2 \gamma})$ is the maximal interval of existence.
Note that $u_0= \varphi_0$, 
and according to Theorem \ref{thm:log-log-phi}, 
the condition  \eqref{eq:negative-initial-energy} suggests that
 $\varphi$ is a blowup solution.
Recall  from \eqref{eq:phi-to-u}, for $T\in (0,\frac{\pi}{4\ga}]$
	\begin{align*}
	u(t, x)
	= \frac{1}{\cos^\frac{n}{2}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	\varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right).
	\end{align*}
Then for all $t \in [0, T)$,
we have
	\begin{align}\label{eq:Du-I1-I2}
	\begin{split}
	\nabla_x u(t,x)
	& = -i \gamma x \frac{\sin(2 \gamma t)}{\cos^{\frac{n}{2}+1}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	\varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \\
	& \quad + \frac{1}{\cos^{\frac{n}{2}+1}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	(e^{tM})^T \nabla \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, \frac{e^{tM} x}{\cos(2 \gamma t)} \right) \\
	& := I_1 + I_2.
	\end{split}
	\end{align}

For $ I_1 $, a change of variable gives
	\begin{align*}
	\vert I_1 \vert_{2}
	= \gamma \sin(2 \gamma t) \left\vert x \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, x \right) \right\vert_{2}.
	\end{align*}
Let $J(t) := \vert x \varphi(t, x) \vert_{2}^2$.
Then \begin{align*}
	J(t)
	= J(0) + J'(0) t + \int_0^t J''(\tau) \,(t - \tau) d\tau.
	\end{align*}
Note that
	\begin{align*}
	|J(0)|
	= \vert x \varphi_0 \vert_2^2
	\leq \vert \varphi_0 \vert_{\mathscr{H}^1}^2,
	\end{align*}
and by Lemma \ref{lem:virial-identity},
we have
\begin{align*}
	|J'(0)|
	& = \left| 4 \Im \int x \overline{\varphi_0} \cdot \nabla \varphi_0 dx \right|
	\leq 4 \| x \varphi_0 \|_2 \| \nabla \varphi_0 \|_2
	\leq  \| \varphi_0 \|_{\mathscr{H}^1}^2, \\
	J''(t)
	& = 8 \mathcal{E}(\varphi_0).
	\end{align*}
Thus
	\begin{align*}
	& \left\vert x \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, x \right) \right\vert_{2}^2
	= J \left( \frac{\tan(2 \gamma t)}{2 \gamma} \right) \\
	& \leq |J(0)|
	+ |J'(0)| \frac{\tan(2 \gamma t)}{2 \gamma}
	+ 4 \mathcal{E}(\varphi_0) \left( \frac{\tan(2 \gamma t)}{2 \gamma} \right)^2 \\
	& \leq \vert \varphi_0 \vert_{\mathscr{H}^1}^2
	+  \vert\varphi_0 \vert_{\mathscr{H}^1}^2 \frac{\tan(2 \gamma T)}{2 \gamma}
	+ 4 \mathcal{E}(\varphi_0) \left( \frac{\tan(2 \gamma T)}{2 \gamma} \right)^2,
	\end{align*}
and so
	\begin{align}\label{eq:I1-estimate}
	\vert I_1 \vert_{2} \leq C(\varphi_0, T).
	\end{align}
	
For $I_2 $, in view of Lemma \ref{lem:e^M}, $e^{tM^T}\in SO(n)$,  
a change of variable gives for $t\in [0,T)$, ($T\le\pi/4\ga$)
	\begin{align*}
	\vert I_2 \vert_{2}
	= \frac{1}{\cos (2 \gamma t)} \left\vert \nabla \varphi\left( \frac{\tan(2 \gamma t)}{2 \gamma}, x \right) \right\vert_{2}.
	\end{align*}
As $t \rightarrow T$,
$\frac{\tan(2 \gamma t)}{2 \gamma} \rightarrow \frac{\tan(2 \gamma T)}{2 \gamma}$,
so by Theorem \ref{thm:log-log-phi},
	\begin{align*}
	\lim_{t \rightarrow T}
	\frac{\left\vert \nabla \varphi \left( \frac{\tan(2 \gamma t)}{2 \gamma}, x \right) \right\vert_{2}}{\vert\nabla Q \vert_{2}}
	\sqrt{ \frac{\frac{\tan(2 \gamma T)}{2 \gamma} - \frac{\tan(2 \gamma t)}{2 \gamma}}
	{\log \left| \log \left( \frac{\tan(2 \gamma T)}{2 \gamma} - \frac{\tan(2 \gamma t)}{2 \gamma} \right) \right|} }
	= \frac{1}{\sqrt{2 \pi}}.
	\end{align*}
Note that as $t \rightarrow T$,
there are
	\begin{align*}
	\frac{\sin(2 \gamma (T - t))}{2 \gamma (T - t)} \rightarrow 1 \qquad \text{and} \qquad
	\frac{\log \sin(2 \gamma (T - t))}{\log(T - t)} \rightarrow 1,
	\end{align*}
so the above blowup rate can be simplified as
	\begin{align*}
	\lim_{t \rightarrow T}
	\frac{\left\vert \nabla \varphi \left( \frac{\tan(2 \gamma t)}{2 \gamma}, x \right) \right\vert_{2}}{\vert \nabla Q \vert_{2}}
	\sqrt{ \frac{T - t}{\log \left| \log (T - t) \right|} }
	= \frac{\cos(2 \gamma T)}{\sqrt{2 \pi}}
	\end{align*}
to yield 
	\begin{align}\label{eq:I2-estimate}
	\lim_{t \rightarrow T}
	\frac{\left\vert I_2 \right\vert_{2}}{\vert \nabla Q\vert_2}
	\sqrt{ \frac{T - t}{\log \left| \log (T - t) \right|} }
	= \frac{1}{\sqrt{2 \pi}}.
	\end{align}
Therefore, combining \eqref{eq:I1-estimate} and \eqref{eq:I2-estimate} we obtain
	\begin{align*}
	\lim_{t \rightarrow T}
	\frac{\left\vert \nabla u \right\vert_{2}}{\vert \nabla Q \vert_{2}}
	\sqrt{ \frac{T - t}{\log \left| \log (T - t) \right|} }
	= \frac{1}{\sqrt{2 \pi}}.
	\end{align*}
\end{proof}

\subsection{Blowup rate at the ground state $Q$}  We conclude with some discussions on the wave collapse rates for \eqref{eq:nls-u}
when the initial data is near $Q$, which could be a subtle issue.  
Notice that this $Q$ is not the ground state for \eqref{eq:nls-u}, instead, it is the one for \eqref{eq:nls-phi}.
 If $\vert u_0 \vert_{2} = \vert Q \vert_{2}$, 
then the wave collapse for \eqref{eq:nls-u}  is different than in the case  $\norm{u_0}_2 >\norm{Q}_2$.  
Applying the transform \eqref{eq:phi-to-u} to the solitary wave $\vphi=e^{it}Q$ 
we can construct a blowup solution with blowup rate $(T-t)\inv$:
	\begin{align}\label{eU:lens-R}
	u(t, x)
	= \frac{1}{\cos^\frac{n}{2}(2 \gamma t)}
	e^{-i \frac{\gamma}{2} |x|^2 \tan(2 \gamma t)}
	e^{i \frac{\tan(2\gamma t)}{2 \gamma}}
	Q\left( \frac{ e^{tM} x}{\cos(2\gamma t)} \right).
	\end{align}
 One easily checks that $u$ blows up at $T=\frac{\pi}{4 \gamma}$ satisfying
	\begin{align*}
	\vert \nabla u \vert_{2}
	\approx \frac{1}{ (\frac{\pi}{2} - 2\ga t)} \vert\nabla Q \vert_{2} \qquad \text{as} \ t \rightarrow T=\frac{\pi}{4 \gamma}.
	\end{align*}  
Note that solutions of the form  \eqref{eU:lens-R} with such blowup time and singularity can also be obtained with  other nonpositive, non-radial bound states $Q_b$ as the profile in place of $Q$,  
where $\norm{Q_b}_2>\norm{Q}_2$ and $n\ge 2$. 
Suppose that $u$ is a blowup solution to \eqref{eq:nls-u} with blowup time $T^* < \frac{\pi}{4 \gamma}$ and $\vert u_0 \vert_{2} = \vert Q \vert_{2}$.
Then by means of \eqref{eq:u-to-phi} we may define a blowup solution  to \eqref{eq:nls-phi} with the same initial data which  blows up at  $T_0 = \frac{\tan(2\gamma T)}{2 \gamma}$.
Merle \cite{Mer} showed that up to the scaling and phase invariances of \eqref{eq:nls-phi},
the only minimal mass blowup solutions are of the form
	\begin{align*}
	\vphi(t,x)=\frac{1}{(T_0 - t)^\frac{n}{2}} e^{-\frac{i |x|^2}{4 (T_0 - t)}} e^\frac{i}{T_0 - t} Q\left( \frac{x}{T_0 - t} - x_0 \right)
	\end{align*}
for some $x_0 \in \mathbb{R}^n$. 
By the $\mathcal{R}$-transform \eqref{eq:phi-to-u} and the uniqueness of \eqref{eq:nls-u} (Proposition \ref{p:blup-altern-conserv}),  
we then establish a characterization for all minimal mass blowup solutions of  \eqref{eq:nls-u}. 
 \begin{proposition}\label{c:charact-mini-pc} Let $\vert u_0 \vert_{2} = \vert Q \vert_{2}$. Let 
 $u$ be a blowup solution of \eqref{eq:nls-u} on $[0,T)$ with $T:=T^* < \frac{\pi}{4 \gamma}$. Then $u$ must assume the following form
 (up to scaling and phase invariance): There exists $x_1\in\R^n$ such that 
\begin{equation}
\begin{aligned}\label{eU:pc-M}
u(t,x)=& \left(\frac{ 2 \gamma \cos(2\gamma T) }{\sin (2 \gamma (T - t))}  \right)^{\frac{n}{2}}
	e^{-i \frac{\gamma}{2} |x|^2 \cot(2 \gamma (T - t))}
	e^{i 2 \gamma \frac{\cos(2 \gamma T) \cos(2 \gamma t)}{\sin(2 \gamma (T - t))}}\\
	&\times Q\left( \frac{2 \gamma \cos(2 \gamma T)  e^{tM}x}{\sin(2 \gamma (T - t))} - x_1 \right).
	\end{aligned}
\end{equation}
Moreover, $\norm{\nabla u}_2\approx (T-t)\inv$ as $t\to T$. 
\end{proposition}   
	Note that the blowup solution given in \eqref{eU:lens-R}  is not covered by  \eqref{eU:pc-M}.
	For example, if $T^*=\pi/8\ga$, $x_1=0$, then \mbox{$u_0(x)=(2\ga)^{n/2} e^{-i\frac{\ga}{2}|x|^2} e^{i2\ga} Q(2\ga x)$.}
Rather,  \eqref{eU:lens-R} can be viewed as a bordering case 
of the assertion in Proposition \ref{c:charact-mini-pc} corresponding to $T^*=\pi/4\ga$. 

\begin{remark}
If the initial value is of the form $u_0 = (1 + \varepsilon) Q$ with $0 < \varepsilon <\sqrt{1+ \frac{\alpha^*}{\| Q \|_2^2}}-1$,
then the corresponding solution to \eqref{eq:nls-u} will blowup at the rate stated in Theorem \ref{thm:log-log-u}.
Indeed,  by the range for $\varepsilon$ and 
the Pohozaev identity $\int |\nabla Q|^2 =\frac{n}{n+2} \int |Q|^{2+\frac{4}{n}} $, 
it is easy to verify  $u_0\in B_{\al^*}$ and condition \eqref{eq:negative-initial-energy}.
\end{remark}

\begin{remark}  
For large initial data one can also derive a general lower bound for the collapse rate. 
If the solution of \eqref{eq:nls-u} satisfies $\lim_{t \rightarrow T^*} \vert\nabla u \vert_{2} = \infty$, 
then there exists  $C=C_{p,n}> 0$ such that
	\begin{align*}
	\vert\nabla u(t) \vert_{2}
	\geq C (T^* - t)^{-(\frac{1}{p-1} - \frac{n-2}{4})}.
	\end{align*}
	 This follows from quite standard argument as in \cite{CazW90} that is used to show the l.w.p 
  and blowup alternative for  the Cauchy problem \eqref{eq:nls-u} on $[0,T^*)$.  
In the $L^2$-critical case $p = 1 + \frac{4}{n}$,  the lower bound becomes
	\begin{align*}
	\vert \nabla u(t) \vert_{2}\geq C (T^* - t)^{-\frac{1}{2}}.
	\end{align*}
\end{remark}


\vspace{.1520in}
\nd{\bf Acknowledgment} The authors would like to thank Remi Carles  and Chenjie Fan for 
 helpful comments and communications.

\end{document}